OpenStudy (anonymous):

ac-5ad-5bc+25bd... How do you factor this expression? :)

OpenStudy (anonymous):

a(c-5d)-5b(c-5d)= (c-5d)(a-5b)

OpenStudy (anonymous):

Thanks :) <3

OpenStudy (saranya):

ac - 5ad - 5bc + 25bd = a(c -5d) -5b(c - 5d) = (a -5b)(c-5d)